OpenStudy (anonymous):

(6x10^6) - (3.1x10^7)

OpenStudy (anonymous):

6,000,000 - 31,000,000 = -25,000,000 = -2.5 x 10 ^7 or 6x10^6 = .6 x 10^7 so .6x10^7 - 3.1x10^7 =